A steel ball of mass 0.50 kg, moving with a velocity 2.0 m/s, strikes a second ball of mass 0.30 kg, initially at rest. The collision is a glancing one, causing the first ball to be deflected by an angle of 30°, with a speed of 1.50 m/s. Determine the velocity of the second ball after the collision, giving both its speed and direction. (6 marks)


1
Expert's answer
2022-01-17T09:59:45-0500

By conservation of momentum:


"Ox:Mv_{xi}=mu_{x}+Mv_f\cos\theta→u_x=0.35\text{ m/s},\\
Oy: 0=mu_y+Mv_f\sin\theta→u_y=-0.375\text{ m/s}.\\\space\\
u=\sqrt{u_x^2+u_y^2}=0.513\text{ m/s},\\\space\\
\theta_u=\arctan\dfrac{u_y}{u_x}=-47°."